# Time series forecasting of multiphase microstructure evolution using deep learning

### v1

@@ Abstract
Microstructure evolution, which plays a critical role in determining materials properties, is commonly simulated by the high-fidelity but computationally expensive phase-field method. To address this, we approximate microstructure evolution as a time series forecasting problem within the domain of deep learning. Our approach involves implementing a cost-effective surrogate model that accurately predicts the spatiotemporal evolution of microstructures, taking an example of spinodal decomposition in binary and ternary mixtures. Our surrogate model combines a convolutional autoencoder to reduce the dimensional representation of these microstructures with convolutional recurrent neural networks to forecast their temporal evolution. We use different variants of recurrent neural networks to compare their efficacy in developing surrogate models for phase-field predictions. On average, our deep learning framework demonstrates excellent accuracy and speedup relative to phase-field simulations. Moreover, we use quantitative measures to demonstrate how surrogate model predictions can replace phase-field timesteps without compromising accuracy in the long-term evolution trajectory. Finally, by emulating a transfer learning approach, our framework performs satisfactorily in predicting new microstructures resulting from alloy composition and physics unknown to the model. As such, our approach demonstrates a useful data-driven alternative or accelerator to microstructure simulation workflow.
I Introduction
@@
Microstructure evolution during materials processing determines the macroscopic properties of the material. Understanding and predicting the time history of microstructure evolution is therefore essential. The phase-field (PF) model is the most high-fidelity method for simulating moving boundary problems such as microstructure evolution [1, 2, 3, 4]. In this method, the diffuse interface representation of phase and grain boundaries obviates the need for explicit interface tracking algorithms. The microstructure evolution relies on the solution of a coupled system of nonlinear partial differential equations (PDEs) involving continuous field variables (e.g., composition, phase, temperature) that evolve both in space and time, making the PF model computationally intensive. Several approaches are employed in the literature to accelerate the PF simulations, including high-performance parallel computing and advanced numerical schemes such as adaptive mesh refinement and nonlinear solution approaches to multiphysics applications [5, 6, 7]. However, the associated computational cost remains high, especially for (i) interfacial dynamics problems, where resolving intricate spatial length scales can be essential as in additive manufacturing [8]; (ii) large-scale long-time simulations [9, 10]; (iii) simulation of multi-component multi-phase material systems [9, 11]; and (iv) when a large number of simulations are required to understand parameter-microstructure relationships [12].
@@
Deep learning (DL) models can provide a viable cost-minimal alternative to microstructure simulation (for recent reviews, see [13, 14]). A DL model encompasses multiple connected machine learning components to learn trends in simulation data involving strong dependencies in space (e.g., images) and time (e.g., sequences) without prior knowledge of the underlying physical mechanisms. These advantages lead to the recent “exploding” interest in applications of DL in materials research. Deep neural models, such as convolutional neural network (CNN) and recurrent neural network (RNN), have architectures particularly suited for the problem at hand, noting that microstructural evolution is essentially a time series problem. An additional challenge, however, remains particularly for handling a large number of model parameters and input/output time frames, affecting the efficiency and speedup of the deep model. To address this, we use a convolutional autoencoder to reduce the dimensional description of the input microstructures in the latent space. These “compressed” microstructures subsequently train the convolutional recurrent neural network (CRNN) to predict future microstructure evolution in the autoencoder latent space. The CRNN framework leverages CNN models to extract the spatial features of the microstructure efficiently and utilizes RNN to forecast microstructure evolution. We are especially interested in RNNs due to their internal “memory” units, which enable them to “remember” microstructure evolution history required for time-series forecasting. We compare RNNs such as simple RNN, long short-term memory (LSTM), and gated recurrent unit (GRU) to gauge their efficacy in developing surrogate models for phase-field predictions. In particular, the LSTM and GRU networks make use of memory cells and gating mechanisms to overcome the vanishing and exploding gradient problems that arise during backpropagation through time (BPTT) in the simple RNN model, enabling them to capture long-term temporal dependencies in the datasets. The GRU offers a simpler alternative to LSTM with fewer tensor operations, allowing for faster training. For a comprehensive tutorial overview of these deep models, refer to [13, 15]. Once the CRNN architecture makes a future prediction, the decoder segment in the autoencoder takes the predicted microstructure in the latent space and reconstructs it back to the original dimensions. The unified approach of convolutional autoencoder + RNN makes an efficient surrogate model – a simplified model designed to represent and substitute a computationally expensive, high-fidelity phase-field model.
@@
Deep learning of microstructure evolution described by PDE-based phase-field simulations remains a challenging topic. Several data-driven approaches exist in the literature for learning and predicting spatiotemporal data generated by phase-field simulations of binary spinodal decomposition and grain microstructure evolution. These studies utilize combination of LSTM and autoencoder architecture [16, 17], combination of two-point spatial correlations, principal component analysis (PCA), and LSTM [18], Eidetic 3D LSTM (by applying convolution to both the spatial and the temporal dimensions) [19], predictive RNN (by modifying the memory cells to accommodate spatiotemporal flow) [20], combination of different dimensional reduction methods and RNN models [21], tensor-decomposition approaches [22, 23], combination of CNN and transfer learning [24], and convolutional artificial neural networks [25].
@@
The majority of the above studies focused on a single deep model with “simple” two-phase material microstructures in which both phases possess the same morphology, which minimizes the inherent variance and randomness during microstructure evolution compared to a three-phase system, where different types of morphology can coexist within the same microstructure [26, 27, 28]. We use different variants of RNN to compare their performance by forecasting the evolution of two-phase and three-phase microstructures. We generate the evolution datasets from phase-field simulations of spinodal decomposition in binary and ternary systems widely found in measurements of metal alloys and polymer blends [28, 29]. The resulting microstructure evolution involves complex curvature-driven interface migration and long-range nonlinear diffusion of multiple chemical species as described by the phase-field equations [30, 31, 32]. To learn a low-dimensional representation of such data on a linear manifold, for example by PCA, may lead to additional spatiotemporal information loss compared to nonlinear embedding techniques such as the autoencoder.
As we show later, the time-evolved future frames predicted by the surrogate model are a good approximation to the corresponding phase-field-generated microstructures. Thus, replacing $n$ number of phase-field timesteps with $n$ number of surrogate model predictions can expedite the overall time evolution of the phase-field workflow. For example, the $n$th frame predicted by the surrogate model can be input to the phase-field solver as the initial configuration for simulating subsequent microstructure evolution. This approach reduces $n$ number of phase-field solve iterations from the solver without any appreciable loss in evolution trajectory. Such a hybrid approach of alternating between phase-field and surrogate models can save significant computing time and storage relative to the full-scale phase-field simulation [18, 16]. However, the question remains on how many phase-field steps can be replaced by the surrogate model to reproduce quantitative microstructure evolution trajectory without loss in accuracy. Moreover, for a comprehensive understanding of the model behavior, it would be interesting to evaluate the predictive performance of the model on new microstructures generated using different material compositions or under different physical mechanisms. Based on the above points, we formulate the following key novel questions that, to our knowledge, have never been addressed before:
  • How does the surrogate model compare for future predictions of two-phase vs. three-phase microstructures?
  • How can we quantitatively determine the phase-field simulation steps that can be “leaped in time” by deep model predictions?
  • How does our surrogate model perform in predicting the evolution trajectory in unseen ternary microstructures?
@@
The structure of the rest of the article is as follows. In Sec. II, we discuss the phase-field models for microstructure evolution. In Sec. III, we briefly describe the components of the deep model employed for phase-field predictions. In Sec. IV, we present the phase-field results as the true microstructures and compare them with the deep model predictions to illustrate the performance of our surrogate model. In Sec. V, we summarize and present a detailed outlook of the present study.
II Phase-field simulations of microstructure evolution: Ground truth
@@
We primarily use a 128 $\times$ 128 domain with a uniform grid spacing of $\Delta x=\Delta y=1.0$, timestep size of $\Delta t=0.005$, and periodic boundary conditions in all directions. In Eq. (1), $\kappa=1$ and $M=1.0$. In Eq. (2), $M_{AA}=M_{BB}=1.0$ and $M_{AB}=0.5$, $\chi_{AB}=\chi_{BC}=\chi_{AC}=3.5$, and $\kappa_{A}=\kappa_{B}=\kappa_{C}=4.0$. We add random noise of variance $\pm$ 0.01 to the initial configuration to initiate phase separation. We implement with the explicit Euler finite-difference algorithm.
We use an in-house phase-field code [30, 31] that generate the spatiotemporal microstructure dataset by solving Eqs. (1) and (2). We collect the microstructure sequences generated after phase-field simulations to construct the dataset for the surrogate model. Each simulation produced 10000 images. The system becomes phase-separated after 500 frames for the ternary mixture and thus is used as a reference time to compare between binary and ternary systems. The time interval between subsequent frames corresponds to 200 timesteps ($\Delta t$ = 0.005).
III Surrogate model architecture: Accelerated phase-field framework
As mentioned in Sec. I, our data-driven approach combines a convolutional autoencoder (encoder-decoder network) with the RNN model. We use simple RNN, LSTM, and GRU as the RNN models. A schematic of our workflow is given in Fig. 1.
@@ III Surrogate model architecture: Accelerated phase-field framework
We use the TensorFlow [36] framework for implementing our data model. We use the standard He [37] initialization to initialize the convolutional kernels (including weights and biases). We use the Adam [38] optimizer to optimize the framework. Unless otherwise noted, we use default configurations as accessible by TensorFlow [36] variables, including the hyperparameter values for both CNN and RNNs. We use the kaggle [39] workspace to implement the surrogate model in Python and run it directly in its notebook environment using GPU accelerator (NVIDIA T4 $\times$ 2 configuration). In order to evaluate the model performance, we use the mean squared error (MSE) as the loss function, which evaluates the error between the ground truth and predictions by [40]:
@@
IV Results and discussion
IV.1 Simulation of ground truth and data preparation for surrogate model
@@
IV.2 One-frame ahead forecasting by RNNs
@@
In Fig. 5, we show short-term and long-term forecasting of microstructure evolution using different RNNs. In the context of the morphology, the short-term predictions closely resemble the ground truth, while significant discrepancies appear in the long-term, which can be visualized in the heat map representing the point-wise difference from the ground truth. Since the differences between the predictions and ground truth accumulate over time series since the training data, visible differences appear in long-term predictions. The forecasting accuracy of different RNN models is shown in Fig. 6. The lowest value of MSE associated with different RNNs signifies that their accuracy increases in the following order: simple RNN < GRU < LSTM. It is evident that LSTM yields better accuracy and long-term prediction ability for the microstructure evolution problem. Also, compared to the binary mixture (MSE < 0.01), the three-phase mixture shows very low MSE (< 0.001). This is likely due to the role of the “third” component in minimizing the symmetry-breaking effect in the ternary mixture, allowing for a much slower, steadier evolution of the microstructure over an extended period of time [30, 31].
@@
IV.3 Multi-frame ahead microstructure forecasting by RNNs
Until now, RNNs are used for temporal extrapolation tests for output one frame forward (Sec. IV.2), i.e., for an input of two consecutive frames ($t-10$ and $t$), predict the next frame in the sequence: $t+10$. How does the RNN model perform when forecasting multiple time frames in the future? In this case, the model’s output is used again as the input for forecasting the next frame. We take the first prediction $t+10$ and include it in the next batch of input clips (i.e., $t$ and $t+10$); then we use the model again to predict the next frame at $t+20$, and we repeat the same process until $t+200$. Using this approach, we ask RNN to output 20 frames (or 200 timesteps) into the future since the last training data at $t=900$. This is illustrated in Fig. 8. The spatial resolution of the output frames progressively decreases when predictions are made increasingly into the future (see data at $t=1010$). The image blurring and information loss in terms of the microstructure spatial features are clearly visible at $t=1010$. The forecasting error is quantified by MSE in Fig. 9. The MSE increases with the predictions increasingly made for the future. This is expected since the point-wise error from each prediction in the time series tends to accumulate, leading to a significant increase in MSE. When we compare different RNNs, their prediction accuracy on average increases in the following order: simple RNN < GRU < LSTM. Let us consider the ternary system. We observe a similar trend earlier in Fig. 6 with low MSE (< 0.001) when the RNNs are asked for one-frame ahead forecasting. However, high MSE (< 0.01) results in this case when forecasting is made for time series in the future based on past predictions instead of the ground truth.
@@ IV.3 Multi-frame ahead microstructure forecasting by RNNs
IV.4 Integration of surrogate and phase-field model frameworks

### v2

@@ Abstract
Microstructure evolution, which plays a critical role in determining materials properties, is commonly simulated by the high-fidelity but computationally expensive phase-field method. To address this, we approximate microstructure evolution as a time series forecasting problem within the domain of deep learning. Our approach involves implementing a cost-effective surrogate model that accurately predicts the spatiotemporal evolution of microstructures, taking an example of spinodal decomposition in binary and ternary mixtures. Our surrogate model combines a convolutional autoencoder to reduce the dimensional representation of these microstructures with convolutional recurrent neural networks to forecast their temporal evolution. We use different variants of recurrent neural networks to compare their efficacy in developing surrogate models for phase-field predictions. On average, our deep learning framework demonstrates excellent accuracy and speedup relative to the “ground truth” phase-field simulations. We use quantitative measures to demonstrate how surrogate model predictions can effectively replace the phase-field timesteps without compromising accuracy in predicting the long-term evolution trajectory. Additionally, by emulating a transfer learning approach, our framework performs satisfactorily in predicting new microstructures resulting from alloy composition and physics unknown to the model. Therefore, our approach offers a useful data-driven alternative and accelerator to the materials microstructure simulation workflow.
keywords:
Deep learning , Autoencoder , CNN , RNN , Microstructure evolution , Phase-field
1 Introduction
@@
Microstructure evolution during materials processing determines the macroscopic properties of the material. Understanding and predicting the time history of microstructure evolution is therefore essential. The phase-field (PF) model is the most high-fidelity method for simulating moving boundary problems such as microstructure evolution [1, 2, 3, 4]. This method uses a diffuse interface representation, which obviates the need for explicit interface tracking. The microstructure evolution relies on the solution of a coupled system of nonlinear partial differential equations (PDEs) involving continuous field variables (e.g., composition, phase, temperature) that evolve both in space and time, making the PF model computationally intensive. Several approaches are employed in the literature to accelerate the PF simulations, including high-performance parallel computing and advanced numerical schemes such as adaptive mesh refinement and nonlinear solution approaches to multiphysics applications [5, 6, 7]. However, the associated computational cost remains high, especially for (i) interfacial dynamics problems, where resolving intricate spatial length scales can be essential as in additive manufacturing [8]; (ii) large-scale long-time simulations [9, 10]; (iii) simulation of multi-component multi-phase material systems [9, 11]; and (iv) when a large number of simulations are required to understand parameter-microstructure relationships [12].
@@
Deep learning (DL) models can provide a viable cost-minimal alternative to microstructure simulation (for recent reviews, see [13, 14]). A DL model encompasses multiple connected machine learning components to learn trends in simulation data involving strong dependencies in space (e.g., images) and time (e.g., sequences) without prior knowledge of the underlying physical mechanisms. These advantages lead to the recent “exploding” interest in applications of DL in materials research. Deep neural models, such as convolutional neural network (CNN) and recurrent neural network (RNN), have architectures particularly suited for the given problem, noting that microstructural evolution is essentially a time series problem. An additional challenge, however, remains particularly for handling a large number of model parameters and input/output time frames, affecting the efficiency and speedup of the DL model. To address this, we use a convolutional autoencoder to reduce the dimensional description of the input microstructures in the latent space. These “compressed” microstructures subsequently train the convolutional recurrent neural network (CRNN) to predict future microstructure evolution in the autoencoder latent space. The CRNN framework leverages CNN models to extract the spatial features of the microstructure efficiently and utilizes RNN to forecast microstructure evolution. We are especially interested in RNNs due to their internal “memory” units, which enable them to “remember” microstructure evolution history required for time-series forecasting. We compare RNNs such as simple RNN, long short-term memory (LSTM), and gated recurrent unit (GRU) to gauge their efficacy in developing surrogate models for phase-field predictions. In particular, the LSTM and GRU networks make use of memory cells and gating mechanisms to overcome the vanishing and exploding gradient problems that arise during backpropagation through time (BPTT) in the simple RNN model, enabling them to capture long-term temporal dependencies in the datasets. The GRU offers a simpler alternative to LSTM with fewer tensor operations, allowing for faster training. For a comprehensive tutorial overview of these deep models, refer to [13, 15]. Once the CRNN architecture makes a future prediction, the decoder segment in the autoencoder takes the predicted microstructure in the latent space and reconstructs it back to the original dimensions. The unified approach of convolutional autoencoder and RNN makes an efficient surrogate model – a simplified model designed to represent and substitute a computationally expensive, high-fidelity phase-field model.
@@
Deep learning of microstructure evolution described by PDE-based phase-field simulations remains a challenging topic. Several data-driven approaches exist in the literature for learning and predicting spatiotemporal data generated by PF simulations of binary spinodal decomposition and grain microstructure evolution. These studies utilize combination of LSTM and autoencoder architecture [16, 17], combination of two-point spatial correlations, principal component analysis (PCA), and LSTM [18], Eidetic 3D LSTM (by applying convolution to both the spatial and the temporal dimensions) [19], predictive RNN (by modifying the memory cells to accommodate spatiotemporal flow) [20], combination of different dimensional reduction methods and RNN models [21], tensor-decomposition approaches [22, 23], combination of CNN and transfer learning [24], and convolutional artificial neural networks [25].
@@
The majority of the above studies focused on a single deep model with “simple” two-phase material microstructures in which both phases possess the same morphology, which minimizes the inherent variance and randomness during microstructure evolution compared to a three-phase system, where different types of morphology can coexist within the same microstructure [26, 27, 28]. We use different variants of RNN to compare their performance by forecasting the evolution of two-phase and three-phase microstructures. We generate the evolution datasets from PF simulations of spinodal decomposition in binary and ternary systems widely found in measurements of metal alloys and polymer blends [28, 29]. The resulting microstructure evolution involves complex curvature-driven interface migration and long-range nonlinear diffusion of multiple chemical species as described by the phase-field equations [30, 31, 32]. To learn a low-dimensional representation of such data on a linear manifold, for example by PCA, may lead to additional spatiotemporal information loss compared to nonlinear embedding techniques such as the autoencoder. Therefore, in this work, we report an autoencoder-based RNN framework as a microstructure learning engine, which can then efficiently be used for the time series forecasting of microstructure evolution without the need to solve the computationally-expensive phase-field governing equations.
@@
The paper is organized as follows. In Sec. 2, we discuss the phase-field models for microstructure evolution. In Sec. 3, we briefly describe the components of the deep model employed for phase-field predictions. In Sec. 4, we present the phase-field results as the true microstructures and compare them with deep model predictions to illustrate the performance of our surrogate model. We discuss the future potential enhancements of deep learning approaches as a microstructure evolution emulator in Sec. 5. We summarize and conclude in Sec. 6.
2 Phase-field simulations of microstructure evolution: Ground truth
@@
We primarily use a 128 $\times$ 128 domain with a uniform grid spacing of $\Delta x=\Delta y=1.0$, timestep size of $\Delta t=0.005$, and periodic boundary conditions in all directions. In Eq. (1), $\kappa=1$ and $M=1.0$. In Eq. (2), $M_{AA}=M_{BB}=1.0$ and $M_{AB}=0.5$, $\chi_{AB}=\chi_{BC}=\chi_{AC}=3.5$, and $\kappa_{A}=\kappa_{B}=\kappa_{C}=4.0$. We add random noise of variance $\pm$ 0.01 to the initial configuration to initiate phase separation. We implement with the semi-implicit Fourier spectral method for numerical integration.
We use an in-house phase-field code [30, 31] that generates the spatiotemporal microstructure dataset by solving Eqs. (1) and (2). We collect the microstructure sequences generated after phase-field simulations to construct the dataset for the surrogate model. Each simulation produced 10000 images. The system becomes phase-separated after 500 frames for the ternary mixture and thus is used as a reference time to compare between binary and ternary systems. The time interval between subsequent frames corresponds to 200 timesteps ($\Delta t$ = 0.005).
3 Surrogate model architecture: Accelerated phase-field framework
As mentioned in Sec. 1, our data-driven approach combines a convolutional autoencoder (encoder-decoder network) with the RNN model. We use simple RNN, LSTM, and GRU as the RNN models. A schematic of our workflow is given in Fig. 1.
@@ 3 Surrogate model architecture: Accelerated phase-field framework
We use the TensorFlow [36] framework for implementing our data model. We use the standard He [37] initialization to initialize the convolutional kernels (including weights and biases). We use the Adam [38] optimizer to optimize the framework. Unless otherwise noted, we use default configurations as accessible by TensorFlow [36] variables, including the hyperparameter values for both CNN and RNNs. We use the Kaggle [39] workspace to implement the surrogate model in Python and run it directly in its notebook environment using graphics processing unit (GPU) accelerator (NVIDIA T4 $\times$ 2 configuration). In order to evaluate the model performance, we use the mean squared error (MSE) as the loss function, which evaluates the error between the ground truth and predictions by [40]:
@@
4 Results and discussion
4.1 Simulation of ground truth and data preparation for surrogate model
@@
4.2 One-frame ahead forecasting by RNNs
@@
In Fig. 6, we show short-term and long-term forecasting of microstructure evolution using different RNNs. In the context of the morphology, the short-term predictions closely resemble the ground truth, while significant discrepancies appear in the long-term, which can be visualized in the heat map representing the point-wise difference from the ground truth. Since the differences between the predictions and ground truth accumulate over time series since the training data, visible differences appear in long-term predictions. The forecasting accuracy of different RNN models is shown in Fig. 7. The lowest value of MSE associated with different RNNs signifies that their accuracy increases in the following order: simple RNN < GRU < LSTM. It is evident that LSTM yields better accuracy and long-term prediction ability for the microstructure evolution problem. Also, compared to the binary mixture (MSE < 0.01), the three-phase mixture shows very low MSE (< 0.001). This is likely due to the role of the “third” component in minimizing the symmetry-breaking effect in the ternary mixture, allowing for a much slower, steadier evolution of the microstructure over an extended period of time [30, 31]. To quantify this, we study the average domain size ($\left\langle R\right\rangle$: a measure of characteristic length scale) of component $A$ in the microstructure by circularly-averaged structure function [42, 43],
where $N$ is the lattice size, $c_{A}(\bm{k},t)$ is the Fourier transform of $c_{A}(\bm{r},t)$, and $k$ is the magnitude of the wave vector $\bm{k}$. Thus, temporal evolution of $\left\langle R\right\rangle$ can be approximated by
$\left\langle R(t)\right\rangle=\frac{\sum S_{A}(k,t)}{\sum kS_{A}(k,t)}.$ (5)
The results are shown in Fig. 5, which demonstrates the larger $\left\langle R(t)\right\rangle$ and different coarsening kinetics of the binary system compared to the ternary counterpart.
Figure 5: Typical evolution of the average domain size ($\left\langle R(t)\right\rangle$) resulting in binary and ternary systems is shown.
@@
4.3 Multi-frame ahead microstructure forecasting by RNNs
Until now, RNNs are used for temporal extrapolation tests for output one frame forward (Sec. 4.2), i.e., for an input of two consecutive frames ($t-10$ and $t$), predict the next frame in the sequence: $t+10$. How does the RNN model perform when forecasting multiple time frames in the future? In this case, the model’s output is used again as the input for forecasting the next frame. We take the first prediction $t+10$ and include it in the next batch of input clips (i.e., $t$ and $t+10$); then we use the model again to predict the next frame at $t+20$; and we repeat the same process until $t+200$. Using this approach, we ask RNN to output 20 frames (or 200 timesteps) into the future since the last training data at $t=900$. This is illustrated in Fig. 9. The spatial resolution of the output frames progressively decreases when predictions are made increasingly into the future (see data at $t=1010$). The image blurring and information loss in terms of the microstructure spatial features are clearly visible at $t=1010$. The forecasting error is quantified by MSE in Fig. 10. The MSE increases with the predictions increasingly made for the future. This is expected since the point-wise error from each prediction in the time series tends to accumulate, leading to a significant increase in MSE. When we compare different RNNs, their prediction accuracy on average increases in the following order: simple RNN < GRU < LSTM. Let us consider the ternary system. We observe a similar trend earlier in Fig. 7 with low MSE (< 0.001) when the RNNs are asked for one-frame ahead forecasting. However, high MSE (< 0.01) results in this case when forecasting is made for time series in the future based on past predictions instead of the ground truth.
@@ 4.3 Multi-frame ahead microstructure forecasting by RNNs
4.4 Integration of surrogate and phase-field model frameworks
As illustrated in Secs. 4.2 and 4.3, the time-evolved future frames predicted by the surrogate model are a good approximation to the corresponding phase-field-generated microstructures. Thus, replacing $n$ number of phase-field timesteps with $n$ number of surrogate model predictions can expedite the overall time evolution of the phase-field workflow. For example, the $n$th frame predicted by the surrogate model can be input to the phase-field solver as the initial configuration for simulating subsequent microstructure evolution. This approach reduces $n$ number of phase-field solve iterations from the solver without any appreciable loss in evolution trajectory. Such a hybrid approach of alternating between phase-field and surrogate models can save significant computing time and storage relative to the full-scale phase-field simulation [18, 16].